### v1

@@ 1 Introduction
The operational efficacy of large language models (LLMs) is governed by the curation and orchestration of their inference-time context [mei2025survey]. As LLM infrastructure evolves from monolithic chat interfaces toward compound agentic systems, context optimization has become both a critical bottleneck and a powerful lever for its enhancement. This establishes the discipline of Context Engineering (CE), which focuses on the principled LLM context optimization to maximize downstream utility and enable continuous self-improvement [hua2025context_engineering_2].
@@
Recent advances demonstrate CE’s efficacy across diverse objectives, such as enhancing vertical-domain performance [zhang2025agentic, cai2025flex, agrawal2025gepa], solving long-horizon tasks [zhang2025recursive_lm, sun2025scaling, hu2025hiagent, kang2025acon], coordinating multiple agents [yuksekgonul2025optimizing, wu2024autogen, ma2026judgeflow], and self-evolution driven by experience [ye2024reevo, lu2026llmasrnnrecurrentlanguagemodel, wei2025evomemory, gao2025survey]. Unlike optimizing LLM parameters, optimizing their context provides distinct methodological advantages: (1) interpretability, by encoding experience in natural language rather than opaque weights; (2) efficiency, by enabling rapid deployment without costly updates of model parameters; (3) modularity, facilitating the composition and transfer of established contexts; and (4) robustness, ensuring immunity to catastrophic forgetting by decoupling capability acquisition from model weights.
@@
However, current CE systems are fundamentally constrained by manually crafted agentic harnesses that impose characteristic inductive biases. At the context representation level, different context structures impose their own trade-offs: case-based trajectories retain rich episodic traces but lack generalization [zhou2025memento]; itemized lists accumulate abstract insights but remain flat and structurally inexpressive [zhang2025agentic]; and graph-based hierarchies offer flexible organization but incur high latency without consistently outperforming naive retrieval [xu2025mem, ai2025memorybench]. At the context optimization level, existing methods exhibit opposing yet equally limiting biases. On one hand, prompt-rewriting approaches such as GEPA [agrawal2025gepa] favor brevity, iteratively refining concise, high-level rules that fail in tasks requiring detailed strategies and deep domain knowledge [zhang2025agentic]. On the other hand, additive-curation approaches [zhang2025agentic, cai2025flex, suzgun2025dynamic, cai2025training] favor verbosity, accumulating context through generation-reflection-curation pipelines that produce noisy context update (due to limited global visibility), incur excessive overhead (due to instance-level optimization), and cause context bloat (due to additive accumulation without holistic synthesis). Ultimately, these heuristic choices restrict CE to a narrow design subspace, precluding the discovery of task-optimal strategies that lie beyond human intuition.
@@
To address these limitations, we introduce Meta Context Engineering (MCE), a framework that supersedes static heuristics through the co-evolution of CE skills and context artifacts. MCE formalizes CE as a bi-level optimization problem, effectively decoupling the engineering strategy (how to represent and optimize context) from the resulting engineered artifact (what context is learned). At the meta-level, we propose agentic skill evolution, in which an agent iteratively refines CE skills [anthropic2025agentskills_blog]: executable instructions and code that govern the CE process. This evolution is driven by agentic crossover, an evolutionary operator that synthesizes superior skills by reasoning across task specifications, historical CE trajectories, and performance metrics. At the base-level, an agent executes these evolved skills to learn from training rollouts and constructs context, in a fully agentic manner. Unlike prior methods that predefine context schemas, our base-agent leverages coding toolkits and file system access to instantiate and optimize context as flexible, programmatic artifacts.
@@
In this manner, MCE replaces heuristic scaffolding with a generic design space. Prior state-of-the-art (SOTA) methods, such as the generation-reflection-curation workflow of agentic context engineering (ACE) [zhang2025agentic], represent singular points within the vast manifold of possible CE skills. By granting AI the full agency and capabilities to generate arbitrary code, invoke other LLMs, and manipulate file structures, MCE can reconstruct such pipelines, discover novel CE architectures, and dynamically adjust CE strategies.
@@
We summarize contributions as follows. ❶ We propose MCE, a bi-level framework that co-evolves CE skills and context artifacts, superseding static CE harnesses with learnable skills. ❷ At the meta-level, we introduce agentic skill evolution, leveraging agent skills [anthropic2025agentskills_blog] as a novel abstraction for evolutionary agent optimization. ❸ At the base-level, we introduce fully agentic context optimization, leveraging coding toolkits and file system access to instantiate context as flexible, programmatic artifacts. ❹ We conduct comprehensive evaluations across five domains, four LLMs, and both offline and online settings. MCE achieves 5.6–53.8% relative improvement over state-of-the-art CE methods (mean of 16.9%), while demonstrating superior context adaptability, transferability, and efficiency in both context usage and training.
@@ 2.1 Agentic Context Engineering
While LLMs possess strong zero-shot capabilities, post-deployment context engineering (CE) is crucial for domain adaptation and self-improvement [khattab2023dspy, hu2025memory, fang2025comprehensive, shinn2023reflexion, yuksekgonul2025optimizing]. Current SOTA CE methods rely on manually designed agentic harnesses: models accumulate experience through exploration-reflection workflows to update contexts with predefined schemas.
@@
However, these harnesses impose characteristic inductive biases. At the context representation level, the trade-offs are distinct: case-based trajectories retain rich episodic traces but lack generalization and abstraction [zhou2025memento, wang2024agent]; itemized lists accumulate abstract insights but remain flat and structurally limited [zhang2025agentic, suzgun2025dynamic]; and hierarchical, graph-based memories incur high latency and complexity without consistently outperforming naive retrieval [chhikara2025mem0, ai2025memorybench, li2025memos, xu2025mem]. At the context optimization level, existing methods exhibit opposing yet equally limiting biases. On one hand, prompt-rewriting approaches such as GEPA [agrawal2025gepa] employ full LLM-based rewrites driven by reflections over sampled trajectories, favoring abstract, high-level rules over detailed domain knowledge (brevity bias). On the other hand, additive-curation approaches such as Dynamic Cheatsheet (DC) [suzgun2025dynamic] and Agentic Context Engineering (ACE) [zhang2025agentic] orchestrate modular agents (generators, reflectors, and curators) to iteratively perform on-policy rollouts, textual reflections, and context updates. These methods utilize localized updates to itemized lists and can additively accumulate insights, but suffer from structural rigidity and context bloat [cai2025flex, cai2025training].
@@
We posit that no single agentic harness is universally optimal, motivating MCE’s dual-level optimization. Our instantiation of MCE is driven by two converging trends. First, agent architectures are shifting from rigid, multi-agent scaffolds toward unified, self-looping, and self-spawning frameworks with maximal agency and minimal yet general tools [anthropic_claude_agent_sdk, manus_app, langchain_deepagents, bolin2026unrolling]. In this architecture, domain specificity can be encapsulated in agent skills: organized instructions, scripts, and resources that agents discover and load dynamically [anthropic2025agentskills_blog]. This motivates our meta-level design: CE as a fully agentic process without restrictive scaffolding, with task specificity injected through learned skills.
@@
Second, coding toolkits and computer (file system) access have became the essential harness of both general and vertical agents [yang2025code, ClaudeCodeDocs2026, manus_app, cheng2026llm]. The prevalence stems largely from the versatility of Turing-complete programming languages, which offer maximal design flexibility [zhang2025darwin, zhang2025aflow], and their inherent verifiability, which facilitates robust training [wei2025_asymmetry_verification, wang2024executable, yang2024swe]. This insight informs our base-level design space: context artifacts consist of files and code, which are general, agent-native representations unconstrained by predefined schemas.
@@ 2.2 Evolutionary Computation with LLMs
Conventional evolutionary computation (EC) relies on manually designed operators, which struggle to capture complex solution structures and domain properties. Recent research demonstrates that LLMs can serve as intelligent genetic operators, leveraging semantic knowledge to generate meaningful variations without explicit rules [wu2024evolutionary].
This paradigm shift enables the exploration of unstructured search spaces and facilitates optimization across varying levels of abstraction [novikov2025alphaevolve, lange2025shinkaevolve]. Prior work in LLM-driven EC has primarily targeted the solution level (e.g., prompts [guo2024connecting], textual solutions [lee2025evolving], numerical parameters [xu2025autoep]), function level (e.g., heuristics [romera2024mathematical, ye2024reevo, liu2024evolution], reward functions [ma2024eureka]), and program level (e.g., search algorithms [novikov2025alphaevolve, hottung2025vrpagent], neural architectures [liu2025alphago], agent workflows [zhang2025evoflow]).
In this work, we introduce agent skills [anthropic2025agentskills_blog] as a novel, integrated level of abstraction for evolutionary optimization. Agent skills are organized folders of instructions, scripts, and resources, that an agent can discover and utilize. We find that this abstraction offers three distinct advantages for meta-level agent optimization. First, by encapsulating instructions, resources, and scripts into a unified representation, skills enable the seamless co-evolution of different solution levels, overcoming the fragmented frameworks often seen in prior co-evolutionary approaches [zhao2025understanding, cen2025beyond, xie2025co, guo2025nested]. Second, skills provide a modular interface for agentic harnesses, decoupling the optimization target from the core agent architecture to enhance stability. Third, this representation facilitates more agentic genetic operators: agents can flexibly inspect and selectively recombine components from ancestor skill directories, enabling more granular and context-aware evolutionary updates.
@@ 3.1 Problem Formulation
Prior CE methods instantiate with predefined components and operators, then optimize directly using fixed procedures. For instance, ACE employs iterative generation-reflection-curation loops to optimize context as an itemized list [zhang2025agentic]. These procedures impose structural biases that may be suboptimal for specific tasks. MCE instead introduces a skill as an executable specification that defines how the context function should be represented and learned from data. Given a skill , a base-level agent executes it to produce a context function . MCE then solves the bi-level problem:
@@ 3.4 Algorithmic Orchestration
Both meta and base-level adopt fully agentic optimizations: each agent interacts with a programming environment through a standard tool set = {Read, Write, Edit, Bash, Glob, Grep, TodoWrite} and produces outputs by manipulating a file system workspace. The read/write permissions of both agents are strictly scoped according to their respective roles and the current iteration. This design is compatible with modern agentic frameworks such as Claude Agent SDK [anthropic_claude_agent_sdk] and LangChain DeepAgents [langchain_deepagents], enabling straightforward integration into existing infrastructure. To facilitate programmatic invocation during rollouts and evaluations, we require the base-agent to implement callable interface(s) with predefined input-output signatures; these interfaces can be task-specific and flexibly defined. They are validated upon completion of each base-agent execution.
@@ 4.1 Experimental Setup
We evaluate MCE across five benchmarks spanning different domains (financial, chemistry, medicine, law, and AI safety) to demonstrate its versatility and generality. Unless otherwise specified, we adhere to the original train/val/test splits but use data subsets due to computational budget constraints. All experiments use the same datasets in the same order during sequential processing to ensure fair comparison across baselines. We include ground-truth labels for CE methods when applicable. Details of the datasets and data processing are provided in Appendix˜A. We briefly describe the datasets below. (1) FiNER (Financial) [loukas2022finer] requires labeling tokens in XBRL financial documents with its entity types. We report the pass@1 prediction accuracy. (2) USPTO-50k (Chemistry) [schneider2016s] requires predicting precursor reactants from product molecules. We report the pass@1 exact match accuracy. (3) Symptom2Disease (Medicine) [gretelai_symptom_to_diagnosis_2023] requires predicting diseases from patient symptom descriptions across 22 disease categories. We report pass@1 prediction accuracy. (4) LawBench (Law) [fei2024lawbench] is a comprehensive Chinese legal benchmark. We use the criminal charge prediction subtask from LawBench, and report micro-F1 score. (5) AEGIS2 (AI Safety) [ghosh2025aegis2] requires classifying user prompts as safe or unsafe and identifying specific violation categories. We report the F1 score.
@@
We compare MCE against the following baselines. To ensure fair comparisons, we adhere to the suggested implementations in ACE [zhang2025agentic], use the identical initial context or prompt templates, and allow a training budget no less than that used in MCE. (1) Base Model: Zero-shot evaluation using the same prompt as MCE but without learned context. (2) In-Context Learning (ICL): Provides task demonstrations in the prompt. Following zhang2025agentic, we include all training samples fitting within the context window. (3) MIPROv2 [opsahl2024optimizing]: We implement the official DSPy implementation with auto="heavy". (4) GEPA [agrawal2025gepa]: We implement the official DSPy implementation with auto="heavy" and match MCE’s total rollout budget. (5) Dynamic Cheatsheet (DC) [suzgun2025dynamic]: We use the official implementation in cumulative mode (DC-CU). (6) ACE [zhang2025agentic]: We use the official implementation with original parameter settings111https://github.com/ace-agent/ace and run offline training for 5 epochs unless playbook accumulation exceeds the context window or validation performance plateaus.
@@
Following zhang2025agentic, we use DeepSeek V3.1 [liu2024deepseek] as the default generator (the model performing inference during training and testing) across all methods and benchmarks. For AEGIS2, safety guardrails necessitate lightweight models; we use Qwen3-8B [qwen3technicalreport] as the generator across all methods. The reflector model is consistently set to DeepSeek V3.1. MCE additionally requires an agentic model; we use MiniMax M2.1 [minimaxai2025minimaxm21] by default. MCE base-agents are allowed to invoke DeepSeek V3.1 during its execution. All models are accessed via OpenRouter [openrouter2025]. In §˜4.3, we show that MiniMax M2.1 does not improve ACE, ruling out knowledge transfer from the agentic model as the source of MCE’s gains.
@@
In experiments, MCE optimizes context for five epochs. We instantiate the MCE agents using the Claude Agent SDK [anthropic_claude_agent_sdk]. To maintain methodological consistency with CE baselines and enable fair comparison, we define the context interface as a one-shot retrieval function: . The system prompts for MCE agents under this instantiation are given in Appendix˜B.
@@
We evaluate all methods under two complementary settings [zhang2025agentic]. (1) Offline: CE methods have access to a training set and iterate over it for multiple epochs to optimize context before evaluation on a held-out test set. (2) Online: CE methods process the test set sequentially, with performance measured only on each instance’s first inference. For MCE, the base-level agent accumulates all processed instances in its file system for continuous learning. We evaluate two MCE variants: one using a fixed skill generated based on task specifications, and another where the base agent operates autonomously without skill guidance.

### v2

@@ 1 Introduction
The operational efficacy of large language models (LLMs) is governed by the curation and orchestration of their inference-time context [Mei et al., 2025]. As LLM infrastructure evolves from monolithic chat interfaces toward compound agentic systems, context optimization has become both a critical bottleneck and a powerful lever for its enhancement. This establishes the discipline of Context Engineering (CE), which focuses on the principled LLM context optimization to maximize downstream utility and enable continuous self-improvement [Hua et al., 2025].
@@
Recent advances demonstrate CE’s efficacy across diverse objectives, such as enhancing vertical-domain performance [Zhang et al., 2026, Cai et al., 2025b, Agrawal et al., 2025], solving long-horizon tasks [Zhang et al., 2025a, Sun et al., 2025, Hu et al., 2025a, Kang et al., 2025], coordinating multiple agents [Yuksekgonul et al., 2025, Wu et al., 2024a, Ma et al., 2026], and self-evolution driven by experience [Ye et al., 2024, Lu et al., 2026, Wei et al., 2025, Gao et al., 2025]. Unlike optimizing LLM parameters, optimizing their context provides distinct methodological advantages: (1) interpretability, by encoding experience in natural language rather than opaque weights; (2) efficiency, by enabling rapid deployment without costly updates of model parameters; (3) modularity, facilitating the composition and transfer of established contexts; and (4) robustness, ensuring immunity to catastrophic forgetting by decoupling capability acquisition from model weights.
@@
However, current CE systems are fundamentally constrained by manually crafted agentic harnesses that impose characteristic inductive biases. At the context representation level, different context structures impose their own trade-offs: case-based trajectories retain rich episodic traces but lack generalization [Zhou et al., 2025]; itemized lists accumulate abstract insights but remain flat and structurally inexpressive [Zhang et al., 2026]; and graph-based hierarchies offer flexible organization but incur high latency without consistently outperforming naive retrieval [Xu et al., 2025a, Ai et al., 2025]. At the context optimization level, existing methods exhibit opposing yet equally limiting biases. On one hand, prompt-rewriting approaches such as GEPA [Agrawal et al., 2025] favor brevity, iteratively refining concise, high-level rules that fail in tasks requiring detailed strategies and deep domain knowledge [Zhang et al., 2026]. On the other hand, additive-curation approaches [Zhang et al., 2026, Cai et al., 2025b, Suzgun et al., 2025, Cai et al., 2025a] favor verbosity, accumulating context through generation-reflection-curation pipelines that produce noisy context update (due to limited global visibility), incur excessive overhead (due to instance-level optimization), and cause context bloat (due to additive accumulation without holistic synthesis). Ultimately, these heuristic choices restrict CE to a narrow design subspace, precluding the discovery of task-optimal strategies that lie beyond human intuition.
@@
To address these limitations, we introduce Meta Context Engineering (MCE), a framework that supersedes static heuristics through the co-evolution of CE skills and context artifacts. MCE formalizes CE as a bi-level optimization problem, effectively decoupling the engineering strategy (how to represent and optimize context) from the resulting engineered artifact (what context is learned). At the meta-level, we propose agentic skill evolution, in which an agent iteratively refines CE skills [Anthropic, 2025]: executable instructions and code that govern the CE process. This evolution is driven by agentic crossover, an evolutionary operator that synthesizes superior skills by reasoning across task specifications, historical CE trajectories, and performance metrics. At the base-level, an agent executes these evolved skills to learn from training rollouts and constructs context, in a fully agentic manner. Unlike prior methods that predefine context schemas, our base-agent leverages coding toolkits and file system access to instantiate and optimize context as flexible, programmatic artifacts.
@@
In this manner, MCE replaces heuristic scaffolding with a generic design space. Prior state-of-the-art (SOTA) methods, such as the generation-reflection-curation workflow of agentic context engineering (ACE) [Zhang et al., 2026], represent singular points within the vast manifold of possible CE skills. By granting AI the full agency and capabilities to generate arbitrary code, invoke other LLMs, and manipulate file structures, MCE can reconstruct such pipelines, discover novel CE architectures, and dynamically adjust CE strategies.
@@
We summarize contributions as follows. ❶ We propose MCE, a bi-level framework that co-evolves CE skills and context artifacts, superseding static CE harnesses with learnable skills. ❷ At the meta-level, we introduce agentic skill evolution, leveraging agent skills [Anthropic, 2025] as a novel abstraction for evolutionary agent optimization. ❸ At the base-level, we introduce fully agentic context optimization, leveraging coding toolkits and file system access to instantiate context as flexible, programmatic artifacts. ❹ We conduct comprehensive evaluations across five domains, four LLMs, and both offline and online settings. MCE achieves 5.6–53.8% relative improvement over state-of-the-art CE methods (mean of 16.9%), while demonstrating superior context adaptability, transferability, and efficiency in both context usage and training.
@@ 2.1 Agentic Context Engineering
While LLMs possess strong zero-shot capabilities, post-deployment context engineering (CE) is crucial for domain adaptation and self-improvement [Khattab et al., 2023, Hu et al., 2025b, Fang et al., 2025, Shinn et al., 2023, Yuksekgonul et al., 2025]. Current SOTA CE methods rely on manually designed agentic harnesses: models accumulate experience through exploration-reflection workflows to update contexts with predefined schemas.
@@
However, these harnesses impose characteristic inductive biases. At the context representation level, the trade-offs are distinct: case-based trajectories retain rich episodic traces but lack generalization and abstraction [Zhou et al., 2025, Wang et al., 2024b]; itemized lists accumulate abstract insights but remain flat and structurally limited [Zhang et al., 2026, Suzgun et al., 2025]; and hierarchical, graph-based memories incur high latency and complexity without consistently outperforming naive retrieval [Chhikara et al., 2025, Ai et al., 2025, Li et al., 2025, Xu et al., 2025a]. At the context optimization level, existing methods exhibit opposing yet equally limiting biases. On one hand, prompt-rewriting approaches such as GEPA [Agrawal et al., 2025] employ full LLM-based rewrites driven by reflections over sampled trajectories, favoring abstract, high-level rules over detailed domain knowledge (brevity bias). On the other hand, additive-curation approaches such as Dynamic Cheatsheet (DC) [Suzgun et al., 2025] and Agentic Context Engineering (ACE) [Zhang et al., 2026] orchestrate modular agents (generators, reflectors, and curators) to iteratively perform on-policy rollouts, textual reflections, and context updates. These methods utilize localized updates to itemized lists and can additively accumulate insights, but suffer from structural rigidity and context bloat [Cai et al., 2025b, a].
@@
We posit that no single agentic harness is universally optimal, motivating MCE’s dual-level optimization. Our instantiation of MCE is driven by two converging trends. First, agent architectures are shifting from rigid, multi-agent scaffolds toward unified, self-looping, and self-spawning frameworks with maximal agency and minimal yet general tools [Anthropic, 2025a, Manus, 2025, langchain-ai, 2025, Bolin, 2026]. In this architecture, domain specificity can be encapsulated in agent skills: organized instructions, scripts, and resources that agents discover and load dynamically [Anthropic, 2025]. This motivates our meta-level design: CE as a fully agentic process without restrictive scaffolding, with task specificity injected through learned skills.
@@
Second, coding toolkits and computer (file system) access have became the essential harness of both general and vertical agents [Yang et al., 2025, Anthropic, 2026, Manus, 2025, Cheng et al., 2026]. The prevalence stems largely from the versatility of Turing-complete programming languages, which offer maximal design flexibility [Zhang et al., 2025c, d], and their inherent verifiability, which facilitates robust training [Wei, 2025, Wang et al., 2024a, Yang et al., 2024]. This insight informs our base-level design space: context artifacts consist of files and code, which are general, agent-native representations unconstrained by predefined schemas.
@@ 2.2 Evolutionary Computation with LLMs
Conventional evolutionary computation (EC) relies on manually designed operators, which struggle to capture complex solution structures and domain properties. Recent research demonstrates that LLMs can serve as intelligent genetic operators, leveraging semantic knowledge to generate meaningful variations without explicit rules [Wu et al., 2024b].
This paradigm shift enables the exploration of unstructured search spaces and facilitates optimization across varying levels of abstraction [Novikov et al., 2025, Lange et al., 2025]. Prior work in LLM-driven EC has primarily targeted the solution level (e.g., prompts [Guo et al., 2024], textual solutions [Lee et al., 2025], numerical parameters [Xu et al., 2025b]), function level (e.g., heuristics [Romera-Paredes et al., 2024, Ye et al., 2024, Liu et al., 2024b], reward functions [Ma et al., 2024]), and program level (e.g., search algorithms [Novikov et al., 2025, Hottung et al., 2025], neural architectures [Liu et al., 2025], agent workflows [Zhang et al., 2025b]).
In this work, we introduce agent skills [Anthropic, 2025] as a novel, integrated level of abstraction for evolutionary optimization. Agent skills are organized folders of instructions, scripts, and resources, that an agent can discover and utilize. We find that this abstraction offers three distinct advantages for meta-level agent optimization. First, by encapsulating instructions, resources, and scripts into a unified representation, skills enable the seamless co-evolution of different solution levels, overcoming the fragmented frameworks often seen in prior co-evolutionary approaches [Zhao et al., 2025, Cen and Tan, 2025, Xie et al., 2025, Guo et al., 2025]. Second, skills provide a modular interface for agentic harnesses, decoupling the optimization target from the core agent architecture to enhance stability. Third, this representation facilitates more agentic genetic operators: agents can flexibly inspect and selectively recombine components from ancestor skill directories, enabling more granular and context-aware evolutionary updates.
@@ 3.1 Problem Formulation
Prior CE methods instantiate with predefined components and operators, then optimize directly using fixed procedures. For instance, ACE employs iterative generation-reflection-curation loops to optimize context as an itemized list [Zhang et al., 2026]. These procedures impose structural biases that may be suboptimal for specific tasks. MCE instead introduces a skill as an executable specification that defines how the context function should be represented and learned from data. Given a skill , a base-level agent executes it to produce a context function . MCE then solves the bi-level problem:
@@ 3.4 Algorithmic Orchestration
Both meta and base-level adopt fully agentic optimizations: each agent interacts with a programming environment through a standard tool set = {Read, Write, Edit, Bash, Glob, Grep, TodoWrite} and produces outputs by manipulating a file system workspace. The read/write permissions of both agents are strictly scoped according to their respective roles and the current iteration. This design is compatible with modern agentic frameworks such as Claude Agent SDK [Anthropic, 2025a] and LangChain DeepAgents [langchain-ai, 2025], enabling straightforward integration into existing infrastructure. To facilitate programmatic invocation during rollouts and evaluations, we require the base-agent to implement callable interface(s) with predefined input-output signatures; these interfaces can be task-specific and flexibly defined. They are validated upon completion of each base-agent execution.
@@ 4.1 Experimental Setup
We evaluate MCE across five benchmarks spanning different domains (financial, chemistry, medicine, law, and AI safety) to demonstrate its versatility and generality. Unless otherwise specified, we adhere to the original train/val/test splits but use data subsets due to computational budget constraints. All experiments use the same datasets in the same order during sequential processing to ensure fair comparison across baselines. We include ground-truth labels for CE methods when applicable. Details of the datasets and data processing are provided in Appendix˜A. We briefly describe the datasets below. (1) FiNER (Financial) [Loukas et al., 2022] requires labeling tokens in XBRL financial documents with its entity types. We report the pass@1 prediction accuracy. (2) USPTO-50k (Chemistry) [Schneider et al., 2016] requires predicting precursor reactants from product molecules. We report the pass@1 exact match accuracy. (3) Symptom2Disease (Medicine) [Gretel AI, 2023] requires predicting diseases from patient symptom descriptions across 22 disease categories. We report pass@1 prediction accuracy. (4) LawBench (Law) [Fei et al., 2024] is a comprehensive Chinese legal benchmark. We use the criminal charge prediction subtask from LawBench, and report micro-F1 score. (5) AEGIS2 (AI Safety) [Ghosh et al., 2025] requires classifying user prompts as safe or unsafe and identifying specific violation categories. We report the F1 score.
@@
We compare MCE against the following baselines. To ensure fair comparisons, we adhere to the suggested implementations in ACE [Zhang et al., 2026], use the identical initial context or prompt templates, and allow a training budget no less than that used in MCE. (1) Base Model: Zero-shot evaluation using the same prompt as MCE but without learned context. (2) In-Context Learning (ICL): Provides task demonstrations in the prompt. Following Zhang et al. [2026], we include all training samples fitting within the context window. (3) MIPROv2 [Opsahl-Ong et al., 2024]: We implement the official DSPy implementation with auto="heavy". (4) GEPA [Agrawal et al., 2025]: We implement the official DSPy implementation with auto="heavy" and match MCE’s total rollout budget. (5) Dynamic Cheatsheet (DC) [Suzgun et al., 2025]: We use the official implementation in cumulative mode (DC-CU). (6) ACE [Zhang et al., 2026]: We use the official implementation with original parameter settings111https://github.com/ace-agent/ace and run offline training for 5 epochs unless playbook accumulation exceeds the context window or validation performance plateaus.
@@
Following Zhang et al. [2026], we use DeepSeek V3.1 [Liu et al., 2024a] as the default generator (the model performing inference during training and testing) across all methods and benchmarks. For AEGIS2, safety guardrails necessitate lightweight models; we use Qwen3-8B [Team, 2025] as the generator across all methods. The reflector model is consistently set to DeepSeek V3.1. MCE additionally requires an agentic model; we use MiniMax M2.1 [MiniMaxAI, 2025] by default. MCE base-agents are allowed to invoke DeepSeek V3.1 during its execution. All models are accessed via OpenRouter [OpenRouter, Inc., 2025]. In §˜4.3, we show that MiniMax M2.1 does not improve ACE, ruling out knowledge transfer from the agentic model as the source of MCE’s gains.
@@
In experiments, MCE optimizes context for five epochs. We instantiate the MCE agents using the Claude Agent SDK [Anthropic, 2025a]. To maintain methodological consistency with CE baselines and enable fair comparison, we define the context interface as a one-shot retrieval function: . The system prompts for MCE agents under this instantiation are given in Appendix˜B.
@@
We evaluate all methods under two complementary settings [Zhang et al., 2026]. (1) Offline: CE methods have access to a training set and iterate over it for multiple epochs to optimize context before evaluation on a held-out test set. (2) Online: CE methods process the test set sequentially, with performance measured only on each instance’s first inference. For MCE, the base-level agent accumulates all processed instances in its file system for continuous learning. We evaluate two MCE variants: one using a fixed skill generated based on task specifications, and another where the base agent operates autonomously without skill guidance.